OpenStudy (anonymous):

An object has a mass of 120 kg on the moon. What is the force of gravity acting on the object on the moon?

OpenStudy (asbhardwaj18):

F=ma as you gave m=120 kg(m is always constant) a here is accn due to gravity which 1/6th that of earth i.e g = 9.8 m/s divided by 6 ..... = 1.6 m/s therefore weight on moon is equal to mg mg=120*1.6 = 192 N